OpenStudy (foxythepirate0984):

2/3 cups of milk are required for 6 pancakes. a. How much milk is needed to make 1 pancake?

OpenStudy (alexandervonhumboldt2):

6=>2/3 1=>x x/1=(2/3)/6 x=2/3*1/6=2/18=simplify

OpenStudy (alexandervonhumboldt2):

make a ratio

OpenStudy (foxythepirate0984):

i dont know how to

OpenStudy (foxythepirate0984):

@YoungStudier Could you help?

OpenStudy (youngstudier):

I don't know..... I always suck at these, sorry :(

OpenStudy (foxythepirate0984):

Its ok

OpenStudy (youngstudier):

@DarrenMadx

OpenStudy (darrenmadx):

sorry i also suck at these too

OpenStudy (foxythepirate0984):

*internal screaming* uh its ok

OpenStudy (youngstudier):

@imqwerty @mitchal